use the equation 2(2-y)=3x what are the values of y, when x= -4, 0, 2/3, and 3?
Found 2 solutions by Alan3354, checkley77:
Answer by Alan3354(31487)

use the equation 2(2-y)=3x what are the values of y, when x= -4, 0, 2/3, and 3?
--------------
2(2-y)=3x
Solve for y
4 - 2y = 3x
-2y = 3x-4
y = 2 - (3x/2)
------------
at x = -4:
y = 2 -(-12/2)
y = 2+6
y = 8
-----------
at x = 0
y = 2
----------
at x = 2/3
y = 2 - 1
y = 1
------------
at x = 3
y = 2 - 9/2
y = -5/2

Answer by checkley77(12569)

2(2-y)=3x x=-4,0, 2/3, & 3.
4-2y=3*-4
-2y=-12-4
-2y=-16
y=-16/-2
y=8 ans.
----------
4-2y=3*0
-2y=-4
y=-4/-2
y=2 ans.
-----------------
4-2y=3*2/3
-2y=6/3-4
-2y=2-4
-2y=-2
y=-2/-2
y=1
------------------
4-2y=3*4
-2y=12-4
-2y=8
y=8/-2
y=-4 ans.